Letv1 = (0,3,6,0),v2 = (0,2,4,6), and v3 = (1,−1,−2,1). Express (4,−1,−2,−11) as a


linear combination of v1,v2, and v3

1
Expert's answer
2022-04-11T07:42:54-0400

To find a linear combination we should solve a system

"\begin{bmatrix}
 0 & 0 & 1 \\
 3 & 2 & -1 \\
 6 & 4 & -2 \\
 0 & 6 & 1
\end{bmatrix} \cdot 
\begin{bmatrix}
 x1 \\
 x2 \\
 x3
\end{bmatrix} =
\begin{bmatrix}
 4 \\
 -1 \\
 -2 \\
 -11
\end{bmatrix}"

That is equivalent to the system of four equations:

"\begin{cases}
 x3 = 4 \\
 3x1+2x2-x3 = -1 \\
6x1 +4x2 -2x3 = -2 \\
6x2+x3 = -11
\end{cases}"

From the first equation get x3 = 4. Substitute this value into the fourth equation obtain 6 x2 = -11 -4, or x2 = -5/2. And from the second equation 3 x1 = -1 +5 +4, so x1 = 8/3. Substituting these values into the third equation, we make sure that it is also fulfilled.

So (4,−1,−2,−11) = 8*v1 -5/2*v + 4*v3